OpenStudy (somegirl1999):

Which number is an irrational number? (1 point) square root of one hundrend twenty-one square root of one hundred fifty square root of four hundred square root of one million

OpenStudy (kamille):

What is irrational number?

OpenStudy (kamille):

Can you give a definition?

OpenStudy (somegirl1999):

sure A Rational Number can be written as a Ratio of two integers (ie a simple fraction).

OpenStudy (kamille):

and does that mean that irrational number can't be written as as ratio of two integers?:)

OpenStudy (somegirl1999):

yes lol srry wrong one hahaha

OpenStudy (ash2326):

@SomeGirl1999 You have given right answer

OpenStudy (somegirl1999):

i have ?

OpenStudy (ash2326):

Try to find square root of each one, 3 are perfect square and 1 is not the one which is not, is your answer

OpenStudy (ash2326):

Yes, you gave the right definition

OpenStudy (somegirl1999):

oh ok

OpenStudy (ash2326):

Are you trying?

OpenStudy (somegirl1999):

im trying but i dont know how

OpenStudy (ash2326):

\[\sqrt{121}=\sqrt{11\times 11}=11\] Do you get this?

OpenStudy (somegirl1999):

yes thanks for the help

OpenStudy (ash2326):

now you understand which is irrational?

OpenStudy (somegirl1999):

yup thanx